# FMDA-OT: Federated Multi-source Domain Adaptation Through Optimal Transport

### v1

@@ 4.1 Datasets
We conducted our experiments using established and widely recognized benchmark datasets, specifically Office-Caltech-10 and VLSC.
VLSC: Dataset consists of images from four different domains: Caltech101, VOC2007, SUN09, and LabelMe. Each domain contains five classes: birds, dogs, chairs, cars, and humans.
Office-Caltech10: Collection consists of images representing four different domains: Amazon, DSLR, Webcam, and Caltech. Each domain comprises 10 categories (i.e., classes) commonly encountered in daily life, household, and office settings, such as computer, bag, car, bike, etc.
@@ 4.2 Results and discussion
In all experiments, we test each model on each target multiple times, and then rely on the average. Currently, the number of tests is 1000.
Table 2: FMDA-OT results on VLSC
@@
[…]
AVG
CMSD[Liu et al., 2021]
37.30
52.74
34.96
31.02
39.00
DS[Liu et al., 2021]
39.52
51.44
41.87
36.44
42.31
TCA+CMSD[Liu et al., 2021]
65.08
54.63
56.69
80.84
64.31
TCA+WAF[Liu et al., 2021]
66.67
54.81
56.40
80.53
64.60
TCA+WDS[Liu et al., 2021]
65.68
56.71
59.71
81.22
65.83
TCA+WDSC[Liu et al., 2021]
65.82
59.69
61.76
81.00
67.05
FMDA-OT (Ours)
69.0
61.0
67.0
96.0
73.25
Table 3: FMDA-OT results on Caltech-Office-10
Target
Method Name
[…]
AVG
ResNet-101[He et al., 2016]
85.4
88.7
99.1
98.2
92.9
DAN[Long et al., 2019]
89.2
91.6
99.5
99.1
94.8
DCTN[Xu et al., 2018]
90.2
92.7
99.4
99.0
95.3
MCD[Saito et al., 2017]
91.5
92.1
99.5
99.1
95.6
$M^{3}SDA$[Peng et al., 2019]
92.2
94.5
99.5
99.2
96.4
FMDA-OT (Ours)
92.0
95.1
99.4
99.2
96.5
The entire system comprises several parameters to adjust, starting with the number of neighbors in spectral clustering, which needs adjustment to enhance clustering quality. We can rely on metrics like the Silhouette index or other internal indices for this evaluation. The regularization parameters in optimal transport, either the Entropy-Regularization or the Class-Regularization, must be adjusted for each source (recommended), or they can have the same values across all sources. However, certain regularization values may lead to mathematical issues, such as division by zero. For our experiments, we provide recommendations regarding the regularization values for each source to enhance the new representation obtained by optimal transport. These parameters can be adjusted locally in each source using the pseudo-labeled data sent from the server (test-driven approach). Generally, it is preferable to minimize the Entropy-Regularization and maximize the Class-Regularization without encountering mathematical errors. In each source, minimizing the Entropy-Regularization below a certain threshold may result in mathematical operational errors. Similarly, maximizing the gap between the Entropy-Regularization and the Class-Regularization increases the likelihood of mathematical operational errors. These mathematical errors are beyond our expertise, as they are inherent to the optimal transport algorithm.
In order to assess the performance of our approaches, we use the Friedman test and Nemenyi test recommended in [Friedman, 1937] using the $Autorank$ package [Herbold, 2020]. The Friedman test is conducted to test the null-hypothesis that all approaches are equivalent in respect of accuracy. If the null hypothesis is rejected, then the Nemenyi test will be performed. In addition, if the average ranks of two approaches differ by at least the critical difference(CD), then it can be concluded that their performances are significantly different. In the Friedman test, we set the significant level ${\alpha=0.05}.$ The figure 5 shows a critical diagram representing a projection of average ranks approaches on enumerated axis. The approaches are ordered from left (the worst) to right (the best) and a thick line which connects the approaches where the average ranks not significantly different (for the level of 5% significance). As shown in figure 6, FMDA-OT achieves significant improvement over the other proposed techniques since during federated learning phase it is stable and the process stops the collaboration for some views when their local quality stars to decrease, which prevents common issue of federaed approaches. Compared to the most cited approaches of the state of the art, the positive impact of using the federated learning based on this theory is:
The statistical analysis was conducted for 7 approaches with 5 paired samples. The family-wise significance level of the tests is alpha=0.050. We rejected the null hypothesis that the population is normal for the approaches CMSD (p=0.000), DS (p=0.000), TCA+CMSD (p=0.000), TCA+WAF (p=0.000), TCA+WDS (p=0.000), TCA+WDSC (p=0.000), and FMDA-OT (p=0.000). Therefore, we assume that not all approaches are normal. Because we have more than two approaches and some of them are not normal, we use the non-parametric Friedman test as omnibus test to determine if there are any significant differences between the median values of the approaches. We use the post-hoc Nemenyi test to infer which differences are significant.
@@
We report the median (MD), the median absolute deviation (MAD) and the mean rank (MR) among all approaches over the samples. Differences between approaches are significant, if the difference of the mean rank is greater than the critical distance CD=1.274 of the Nemenyi test. We reject the null hypothesis (p=0.000) of the Friedman test that there is no difference in the central tendency of the approaches CMSD (MD=37.300+-2.025, MAD=2.340, MR=6.800), DS (MD=41.870+-1.400, MAD=2.350, MR=6.200), TCA+CMSD (MD=64.310+-4.195, MAD=7.620, MR=4.600), TCA+WAF (MD=64.600+-5.135, MAD=8.200, MR=4.000), TCA+WDS (MD=65.680+-3.060, MAD=5.970, MR=3.000), TCA+WDSC (MD=65.820+-2.645, MAD=4.060, MR=2.400), and FMDA-OT (MD=69.000+-3.125, MAD=4.250, MR=1.000). Therefore, we assume that there is a statistically significant difference between the median values of the approaches. Based on the post-hoc Nemenyi test, we assume that there are no significant differences within the following groups: CMSD and DS; TCA+CMSD and TCA+WAF; TCA+WAF and TCA+WDS; TCA+WDS and TCA+WDSC. All other differences are significant.
@@
In our work, we propose an architecture for privacy preservation and domain adaptation in the field of multi-source domain adaptation, allowing for model training in an unsupervised manner without access to the data sources. This approach is essential in today’s landscape to uphold client privacy across various industries. The architecture consists of two adaptation phases: firstly, employing optimal transport selectively for the sources, and secondly, the server of Federated Learning (FL) guiding the adaptation by defining the importance of each source. Our approach demonstrates its effectiveness on benchmark datasets. There are several exciting directions to explore for further development and privacy preservation. Firstly, we believe that our architecture’s performance can be enhanced by developing or utilizing more sophisticated FL systems. Secondly, the new representation obtained by optimal transport in each source can be replaced by alternative representations from other algorithms that may offer greater discrimination.

### v2

@@ 4.1 Datasets
We conducted our experiments using two well-established benchmark datasets: Office-Caltech-10 and VLSC.
VLSC: This dataset includes images from four different domains: Caltech101, VOC2007, SUN09, and LabelMe. Each domain contains five classes: birds, dogs, chairs, cars, and humans.
Office-Caltech10: This collection consists of images from four domains: Amazon, DSLR, Webcam, and Caltech. Each domain comprises 10 categories commonly encountered in daily life, household, and office settings, such as computers, bags, cars, and bikes.
4.2 Hyperparameters Tuning
In our experiments, the hyperparameters were selected using a test-driven approach to ensure the performance. We also provide recommendations for tuning these parameters in the following discussion, as they can vary depending on the data distribution. Specifically, during our experiments, we set the entropy regularization to 50 and the class regularization to 5000 for the optimal transport method. For the clustering algorithm (spectral clustering), the number of neighbors was set to 12. It’s important to note that these parameters are inherently sensitive to the underlying data distribution, and adjustments may be necessary depending on the specific characteristics of the data used.
@@ 4.3 Results and Discussion
In all experiments, we evaluated each model on each target multiple times, and relied on the average performance. Currently, the number of tests conducted is 1000.
Table 2: CMDA-OT results on VLSC
@@
CMSD [Liu et al., 2021] 37.30 52.74 34.96 31.02 39.00
DS [Liu et al., 2021] 39.52 51.44 41.87 36.44 42.31
TCA+CMSD [Liu et al., 2021] 65.08 54.63 56.69 80.84 64.31
TCA+WAF [Liu et al., 2021] 66.67 54.81 56.40 80.53 64.60
TCA+WDS [Liu et al., 2021] 65.68 56.71 59.71 81.22 65.83
TCA+WDSC [Liu et al., 2021] 65.82 59.69 61.76 81.00 67.05
CMDA-OT (Ours) 69.0 61.0 67.0 96.0 73.25
Table 3: CMDA-OT results on Caltech-Office-10
Target
@@
ResNet-101 [He et al., 2016] 85.4 88.7 99.1 98.2 92.9
DAN [Long et al., 2019] 89.2 91.6 99.5 99.1 94.8
DCTN [Xu et al., 2018] 90.2 92.7 99.4 99.0 95.3
MCD [Saito et al., 2017] 91.5 92.1 99.5 99.1 95.6
$M^{3}SDA$ [Peng et al., 2019] 92.2 94.5 99.5 99.2 96.4
CMDA-OT (Ours) 92.0 95.1 99.4 99.2 96.5
The system includes some parameters that need to be tuned for optimal performance. One key parameter is the number of neighbors in spectral clustering, which needs adjustment to enhance clustering quality. Metrics like the Silhouette index or other internal indices can be used for this evaluation. The regularization parameters in optimal transport, whether Entropy-Regularization or Class-Regularization, must also be adjusted for each source (preferable) or uniformly across all sources. Certain regularization values may lead to mathematical issues, such as division by zero. For our experiments, we provide recommendations regarding the regularization values for each source to enhance the new representation obtained by optimal transport. These parameters can be adjusted locally in each source using pseudo-labeled validation data (test-driven approach). Generally, it is preferable to minimize Entropy-Regularization and maximize Class-Regularization without encountering mathematical errors. Minimizing Entropy-Regularization below a certain threshold or maximizing the gap between Entropy-Regularization and Class-Regularization may result in mathematical operational errors. These mathematical errors are inherent to the optimal transport algorithm and beyond our control.
To assess the performance of our approaches, we use the Friedman test and Nemenyi test as recommended in [Friedman, 1937] using the $Autorank$ package [Herbold, 2020]. The Friedman test is conducted to test the null hypothesis that all approaches are equivalent in terms of accuracy. If the null hypothesis is rejected, the Nemenyi test is performed. If the average ranks of two approaches differ by at least the critical difference (CD), their performances are significantly different. In the Friedman test, we set the significance level at ${\alpha=0.05}$.
Figure 5 shows a critical diagram representing the projection of average ranks of approaches on an enumerated axis. The approaches are ordered from left (worst) to right (best), with a thick line connecting the approaches where the average ranks are not significantly different at a 5% significance level. As shown in Figure 6, CMDA-OT achieves significant improvement over other proposed techniques, demonstrating stability during the federated learning phase. This process stops the collaboration for certain views when their local quality starts to decrease, preventing common issues in federated approaches. Compared to the most cited state-of-the-art approaches, the positive impact of using federated learning based on this theory is evident.
The statistical analysis was conducted for 7 approaches with 5 paired samples. The family-wise significance level of the tests is alpha=0.050. We rejected the null hypothesis that the population is normal for the approaches CMSD (p=0.000), DS (p=0.000), TCA+CMSD (p=0.000), TCA+WAF (p=0.000), TCA+WDS (p=0.000), TCA+WDSC (p=0.000), and CMDA-OT (p=0.000). Therefore, we assume that not all approaches are normal. Because we have more than two approaches and some of them are not normal, we use the non-parametric Friedman test as omnibus test to determine if there are any significant differences between the median values of the approaches. We use the post-hoc Nemenyi test to infer which differences are significant.
@@
We report the median (MD), the median absolute deviation (MAD) and the mean rank (MR) among all approaches over the samples. Differences between approaches are significant, if the difference of the mean rank is greater than the critical distance CD=1.274 of the Nemenyi test. We reject the null hypothesis (p=0.000) of the Friedman test that there is no difference in the central tendency of the approaches CMSD (MD=37.300+-2.025, MAD=2.340, MR=6.800), DS (MD=41.870+-1.400, MAD=2.350, MR=6.200), TCA+CMSD (MD=64.310+-4.195, MAD=7.620, MR=4.600), TCA+WAF (MD=64.600+-5.135, MAD=8.200, MR=4.000), TCA+WDS (MD=65.680+-3.060, MAD=5.970, MR=3.000), TCA+WDSC (MD=65.820+-2.645, MAD=4.060, MR=2.400), and CMDA-OT (MD=69.000+-3.125, MAD=4.250, MR=1.000). Therefore, we assume that there is a statistically significant difference between the median values of the approaches. Based on the post-hoc Nemenyi test, we assume that there are no significant differences within the following groups: CMSD and DS; TCA+CMSD and TCA+WAF; TCA+WAF and TCA+WDS; TCA+WDS and TCA+WDSC. All other differences are significant.
@@
In this work, we propose an architecture designed to ensure privacy preservation and facilitate domain adaptation in multi-source domain adaptation scenarios. Our approach enables unsupervised model training without access to the data sources, which is crucial for maintaining client privacy across diverse industries today. The architecture operates in two adaptation phases: initially, it uses selective optimal transport for each source, and subsequently, the collaborative learning, with Federated Learning (FL) server directs the adaptation by determining the significance of each source. Our method has proven effective on benchmark datasets. Looking ahead, there are several promising avenues for further enhancement and privacy preservation. Firstly, we believe that the performance of our architecture can be improved by developing or integrating more advanced collaborative or FL systems. Secondly, the new representations obtained via optimal transport at each source can be substituted with alternative representations from other algorithms that may offer superior discriminative capabilities.